OpenStudy (moongazer):

What percent profit was made on a pair of pants that was bough for P600 and sold for P630 ?

OpenStudy (anonymous):

30/600 is 1/20th or 5%

OpenStudy (moongazer):

why?

OpenStudy (moongazer):

please give me another exaample and let me solve it

OpenStudy (moongazer):

after you explain :)

OpenStudy (anonymous):

Cost price 100, profit 10

OpenStudy (moongazer):

?

OpenStudy (anonymous):

Cost price 100, profit 10---%profit?

OpenStudy (anonymous):

What does "percent" mean?

OpenStudy (anonymous):

It means "per 100" so if u make 10 per 100 u have made 10%

OpenStudy (anonymous):

Take your original question, I said 30/600 is 1/20th or 5% I can also put 30/600 = 5/100 = 5 percent (5%)

OpenStudy (anonymous):

5% is another way of writing 5/100.

OpenStudy (moongazer):

?

OpenStudy (moongazer):

another example? and i will answer

OpenStudy (anonymous):

200 cost, 18 profit...

OpenStudy (moongazer):

?

OpenStudy (moongazer):

:) give me another example then i will answer for practice

OpenStudy (anonymous):

200 cost, 18 profit

OpenStudy (anonymous):

% profit?

OpenStudy (moongazer):

?

OpenStudy (moongazer):

what is the question?

OpenStudy (moongazer):

?

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

Come on, u can do that in your head..

OpenStudy (moongazer):

?

OpenStudy (moongazer):

9%

OpenStudy (moongazer):

?

OpenStudy (moongazer):

right?

OpenStudy (moongazer):

?

OpenStudy (moongazer):

am i correct?

OpenStudy (anonymous):

Yup. 250, 25 profit...

OpenStudy (moongazer):

?

OpenStudy (moongazer):

10%

OpenStudy (anonymous):

Same, %profit?

OpenStudy (moongazer):

?

OpenStudy (moongazer):

could you ask in a sentence manner?

OpenStudy (anonymous):

Correct. If u made 7% profit and it was 40$, what was the cost price?

OpenStudy (moongazer):

?

OpenStudy (moongazer):

42.8?

OpenStudy (moongazer):

am i right?

OpenStudy (anonymous):

The cost price is 100%